1 A car accelerates from rest at 4 m/s2. What is the velocity of the car after 4 seconds? (Answer: 16 m/s

1

A car accelerates from rest at 4 m/s2. What is the velocity of the car after 4 seconds? (Answer: 16 m/s

Grade:12

1 Answers

Yash Jain
9 years ago
v = u + at ( 1st eq. of motion)
for the case given, u = 0m/s, a = 4m/s2, t = 4s
So, v = 0 + 4*4 = 16m/s